# Towards Efficient Real-Time Video Motion Transfer via Generative Time Series Modeling

### v1

@@ 5.3 Diverse Video Sample Prediction
Figure 14: Qualitative results for VoxCeleb dataset in reconstruction mode in generating diverse samples using GRU-NF (upper panel) and VRNN (lower panel) for keypoint prediction
Figure 15: Qualitative results for VoxCeleb dataset in transfer mode in generating diverse samples using GRU-NF (upper panel) and VRNN (lower panel) for keypoint prediction
5.4 Comparative Analysis of Algorithms
@@
Single Video Sample Prediction: Among the models evaluated—VRNN, VAE, NF, GRU, and GRU-NF, VRNN consistently outperformed the others in the task of direct prediction on video keypoints. This superior performance is aligned with the inherent strengths of the VRNN architecture, which is particularly well-suited for modeling sequential data that exhibits complex temporal structure, high intrinsic variability, and a high signal-to-noise ratio. VRNNs are known to effectively handle time series with abrupt changes or non-smooth dynamics that are not attributable to random noise—an accurate characterization of keypoint trajectories. These trajectories are often non-deterministic yet geometrically constrained, making them difficult to predict using purely deterministic or memory-limited models.
@@
Traditional RNNs, such as GRUs, struggle with capturing such abrupt transitions and complex variability, as their deterministic nature and limited memory can fail to represent latent stochastic dynamics. VAEs, while effective at learning probabilistic latent representations and generating diverse samples, are not inherently designed to model temporal dependencies, and thus may underperform in capturing the evolving structure of keypoint motion. Normalizing Flow (NF) models, despite offering exact likelihood computation and expressiveness in latent space, can struggle to model temporal coherence when applied directly to sequences, as they lack built-in mechanisms to capture sequence dependencies. The hybrid GRU-NF, although combining temporal modeling from GRU with expressiveness from NF, still underperforms compared to VRNN. Since two parts (GRU and NF) of the model aren’t learning collaboratively, which means the GRU isn’t updated based on the latent sample generted by NF, its hidden states may not provide the most useful information for the NF especially in the presence of complex, high-variation motion patterns. Similarly, the NF may focus on fitting the data distribution well, but without proper feedback from the temporal learning component, it might not fully leverage the temporal structure encoded by the GRU. This contrasts with VRNN, where the recurrent structure and the latent variables are trained end-to-end in a tightly coupled fashion, along with minimizing reconstruction loss—ensuring that the model learns both temporal dynamics and stochastic variation in a coordinated way.
@@
Diverse Video Sample Prediction: Except the GRU all the models studied in this work have a latent space that helps them to produce more than one plausible outcome. Although VAE and NF are able to produce diverse samples, they lack the ability to capture temporal diversity, limiting their effectiveness in modeling dynamic, time-evolving structures. Therefore, we have excluded these models when considering diverse video sample prediction.
@@
The results from our simulations using both VRNN and GRU-NF show that each model utilizes its latent space uniquely. Consequently, VRNN demonstrates better performance in terms of MAE, whereas GRU-NF provides better diversity among samples. This distinction suggests different potential applications for each model. VRNN is a VAE based architecture which relies on approximate likelihood estimation, that can hinder its ability to capture the full complexity of the underlying data distribution. Since GRU-NF is based on NF architecture that uses exact likelihood estimation, it learns a more expressive latent space with flexible transformations, leading to a wider and more complex distribution. The substantial differences between the latent samples lead to a comparatively higher deviation from ground truth at the cost of a wider variety of predictions. This trade-off highlights the strength of exact likelihood methods like NF in capturing the complete probability distribution, which is especially valuable for applications requiring probabilistic forecasting or uncertainty quantification.
Table 8: Diversity vs. fidelity results of VoxCeleb dataset on transfer mode
APD MAE SSIM APD to MAE ratio
Mean Median
VRNN 0.0083 0.0174 0.9412 0.4451 0.4871
GRU-NF 0.0336 0.0464 0.7861 0.9256 0.9397
Table 9: Diversity vs. fidelity results of Bair dataset on reconstruction mode
APD MAE SSIM APD to MAE ratio
Mean Median
VRNN 0.0096 0.0787 0.7494 0.3603 0.2977
GRU-NF 0.0269 0.0853 0.7377 1.3224 1.2295
Table 10: Diversity vs. fidelity results of Bair dataset on transfer mode
APD MAE SSIM APD to MAE ratio
Mean Median
VRNN 0.0094 0.0306 0.9186 0.5128 0.4408
GRU-NF 0.0268 0.0393 0.8996 1.2062 1.2283
@@ 6 Conclusions and Future Work
In this research, we aim to investigate the strengths and limitations of generative time series models for both single and diverse sample prediction in video datasets. We explore a combination of the GRU which is a sequence prediction model with two different generative models, VAE and NF, implemented individually. The first combination, known as VRNN, pairs GRU with VAE, while the second, GRU-NF, integrates GRU with Normalizing Flows.
@@
Our findings indicate that in single video sample prediction, VRNN surpasses GRU, VAE, NF and GRU-NF by a significant margin, and can potentially enable 2x additional bandwidth reduction in existing motion transfer pipelines. However, in diverse video sample prediction, GRU-NF proves superior in terms of the diversity-fidelity tradeoff. While VRNN generates samples closely resembling the ground truth with minimal variation among them, GRU-NF produces more diverse samples without significantly compromising visual quality. This highlights the fundamental distinction between the two approaches:VRNN prioritizes accuracy at the expense of diversity, whereas GRU-NF does the opposite.
@@
The choice between these models depends on the specific application for real-time motion transfer depending on whether the task requires precision or diversity. In scenarios such as video conferencing and remote patient monitoring, where accurate forecasting is critical, VRNN is the preferred choice. Conversely, applications like VR gaming and vision-based anomaly detection in the manufacturing industry require diverse sample generation at future timesteps with high fidelity, making GRU-NF the more suitable option. Our future work will focus on investigation of these tradeoffs involved in using various generative time series models for different realtime motion transfer applications on resource constrained edge computing platforms.
Acknowledgements
@@
Tasmiah Haque was supported for this work by an Intel grant. The authors would like to acknowledge the Pacific Research Platform, NSF Project ACI-1541349, and Larry Smarr (PI, Calit2 at UCSD) for providing the computing infrastructure used in this project. The authors also thank Md Mushfiqur Rahaman for helpful technical discussions and computing support.
@@ A. Motion Transfer Backbone (FOMM)
For motion transfer, we utilize the original implementation of FOMM and retain its default configurations for our experiments. For architectural and implementation details, please refer to the supplementary material of [51].
@@ B.1 VRNN
The VRNN model is based on a single-layer GRU with 256 hidden units that captures temporal dependencies in 60-dimensional keypoint sequences. It integrates latent-variable modeling using a VAE-like structure. At each timestep, a 20 dimensional latent vector is sampled from a Gaussian distribution using prior and posterior networks, both implemented as feedforward layers with 128 hidden units. The latent code is concatenated with the GRU hidden state and passed through a decoder with a 128-unit hidden layer followed by a linear layer mapping to 60-dimensional keypoints. For Mgif and VoxCeleb datasets, it takes 12 or 6 observed keypoint sequences and predicts the next 12 or 6 steps, respectively, and for Bair Dataset, it takes 15 or 5 observed keypoint sequences and predicts the next 15 or 5 steps, respectively.
@@ B.2 GRU-NF
We use a probabilistic sequential generative model that combines a three-layer GRU network with 2048 hidden units per layer, followed by an NF component implemented using 12 RealNVP blocks. The GRU encodes the 60-dimensional input into a hidden representation which is then used to generate a sequence of 60-dimensional latent variables through the RealNVP blocks. Each block contains a pair of scale and translation subnetworks, each composed of five fully connected layers with 2048 hidden units and Tanh/ReLU activations, respectively. These blocks are interleaved with Batch Normalizing layers to stabilize training. For Mgif and VoxCeleb datasets, the model takes an input sequence of 12 or 6 time steps and forecasts the next 12 or 6 timesteps, respectively, and for Bair Dataset, it takes 15 or 5 time steps and predicts the next 15 or 5 steps, respectively.

### v2

@@ 5.3 Diverse Video Sample Prediction
Figure 14: Qualitative results for VoxCeleb dataset in reconstruction mode in generating diverse samples using (a) GRU-NF and (b) VRNN for keypoint prediction
Figure 15: Qualitative results for VoxCeleb dataset in transfer mode in generating diverse samples using (a) GRU-NF and (b) VRNN for keypoint prediction
5.4 Throughput, Compute and Bandwidth for Real-Time Use
Setup. We measure end-to-end latency of the full pipeline of our model: keypoint extraction → forecasting → video generation for single video sample prediction, on an NVIDIA RTX 3090 (24 GB VRAM), CUDA 12.x, PyTorch FP32 (Full precision 32 bit, no quantization) using 256×256 video frames from the VoxCeleb dataset. We compare this versus other keypoint-based motion transfer models; namely the Thin Plate Spline (TPS) model integrated with VRNN and the Neural Talking Head model integrated with VRNN by measuring latency for them on the same hardware.
Inference time and fps estimate. For each model we process 36 consecutive frames arranged as blocks of 6 observed frames and 6 predicted frames in a repeating pattern. Thus, over 36 frames we have 18 observed (used for keypoint extraction) and 18 predicted (produced by the VRNN), while video generation runs on all 36 frames. We record per-stage wall-clock time and aggregate these to obtain the total inference time for 36 frames. To report performance as per standard videoconferencing setups, we use linear scaling to obtain the inference time T for 30 frames; following which the frames per second (fps) is estimated as 30/T.
TOPS estimate. Our steady-state compute accounting follows stride-2 operation across the 3 stages of the pipeline: computations during keypoint extraction and keypoint forecasting occur only during half of the total 30 frames, whereas video generation runs on every frame as mentioned earlier. We incorporate these duty cycle considerations when converting per-stage operation counts to model-required Trillion Operations per second (TOPS). Operation counts are obtained using fvcore’s FlopCountAnalysis on representative inputs; fvcore traces the model and reports a per-module flop (MAC) estimate [fvcore]. We then report TOPS = (FLOPs per output frame × fps) / (u × 10¹²) assuming an average device utilization of 60% [gao2024empirical], i.e. u=0.6. These differences in compute across the different stages are taken into account during our TOPS calculations and the reasoning generalizes to other strides.
[…]
Table 11: Comparison of inference time for the entire pipeline with other motion transfer models. For each criteria i.e. total inference time and total compute the best model is highlighted in bold and the second best is highlighted by underline, lower the better. Measurements were taken on 36 frames and linearly scaled to 30 frames.
Model
Total
time (s)
Total
TOPS@30 fps
Keypoint extraction Keypoint forecast Video generation
time (s)
TOPS
time (s)
TOPS
time (s)
TOPS
FOMM+VRNN 0.8176 2.7194 0.0336 0.0310 0.0743 0.0001 0.7097 2.6883
TPS+VRNN 1.0627 6.4210 0.0283 0.0592 0.0878 0.0001 0.9466 6.3617
Neural Talking
Head+VRNN
1.2999 10.0026 0.0893 0.1958 0.0782 0.0001 1.1323 9.8067
Table 12: Comparison of throughput, compute, bandwidth savings and accuracy with other motion transfer models. For each column the best model is highlighted in bold and the second best is highlighted by underline. For throughput and bandwidth saving, higher is better; for compute, MAE and JEDi, lower is better.
Model
Throughput
Compute
Bandwidth
Saving ($\times$)
MAE JEDi
FOMM+VRNN 36.69 2.7194 17.36 0.0511 0.7079
TPS+VRNN 28.23 6.4210 10.40 0.0499 0.5716
Neural Talking Head+VRNN 23.08 10.0026 20.42 0.0619 0.738
We benchmark all 3 models on VoxCeleb dataset in reconstruction mode using a 6-6 prediction horizon. The inference time and compute per-stage (keypoint extraction, keypoint forecast and video generation) to generate 30 consecutive frames are displayed in Table 12 and the throughput, total compute (TOPS) for the entire pipeline, bandwidth saving, MAE and JEDi are displayed in Table 12 for all three models. From the results we can observe that, FOMM+VRNN is the most suitable choice for real-time, resource-constrained deployment: it takes the lowest time (0.8176 s) to generate consecutive 30 frames which is 23.07% lower than the time required for the second best model TPS+VRNN. Moreover, it is the fastest at 36.69 fps and needs the least compute (2.7194 TOPS) which is 29.97% faster and 57.65% lesser than TPS+VRNN, respectively. The lowest bitrate estimate is 24.48 kbps for Neural Talking Head+VRNN models which provides the highest bandwidth saving 20.42x vs a 500 kbps baseline, however it attains the worst MAE and JEDi among the competing models. FOMM+VRNN model requires 28.8 kbps which leads to 17.36× saving, providing near-best bandwidth reduction while also achieving near-best MAE and JEDi. Therefore, the substantial gains in fps, compute, and bandwidth make FOMM+VRNN the best overall real-time option.
Selection rationale. For real-time, resource-constrained deployment (edge or interactive), FOMM+VRNN offers the best Pareto balance: it is the fastest (36.69 fps), requires the least compute (2.7194 TOPS), and achieves near-best bandwidth reduction, while keeping MAE within $\approx 2\%$ of the most accurate variant. TPS+VRNN remains a quality-first alternative when maximal perceptual accuracy (MAE/JEDi) is prioritized and runtime/compute are less constrained.
6 Discussion
@@
6.1 Single Video Sample Prediction
Among the evaluated models (VRNN, VAE, NF, GRU, and GRU-NF), the results indicate that performance depends on the dataset regime and the prediction horizon, rather than one model dominating uniformly across all settings. This behavior aligns with the different inductive biases of the models and with how conditional uncertainty evolves with horizon. VRNN is particularly well-suited for modeling sequential data that exhibits complex temporal structure, high intrinsic variability, and a high signal-to-noise ratio. It explicitly introduces time-dependent latent variables coupled with recurrence, enabling it to represent latent stochastic factors that influence future motion and become increasingly important as the forecast horizon grows. Consequently, VRNN shows clearer advantages in settings that are more conditionally multi-modal i.e., multiple plausible futures from similar prefixes, which is most evident on BAIR at longer horizons. In contrast, GRU is deterministic and often provides stable rollouts when the conditional dynamics are closer to unimodal e.g., phase-locked or low-amplitude motion. This helps explain why GRU shows superior performance in MGIF and VoxCeleb datasets, particularly at longer horizons where small temporal inconsistencies can accumulate.
@@ 6.1 Single Video Sample Prediction
VAEs, while effective at learning probabilistic latent representations and generating diverse samples, are not inherently designed to model temporal dependencies, and thus may underperform in capturing the evolving structure of keypoint motion. NF models, despite offering exact likelihood computation and expressiveness in latent space, can struggle to model temporal coherence when applied directly to sequences, as they lack built-in mechanisms to capture sequence dependencies. The hybrid GRU-NF, although combining temporal modeling from GRU with expressiveness from NF, still underperforms compared to VRNN. Since two parts (GRU and NF) of the model aren’t learning collaboratively, which means the GRU isn’t updated based on the latent sample generated by NF, its hidden states may not provide the most useful information for the NF especially in the presence of complex, high-variation motion patterns. Similarly, the NF may focus on fitting the data distribution well, but without proper feedback from the temporal learning component, it might not fully leverage the temporal structure encoded by the GRU. This contrasts with VRNN, where the recurrent structure and the latent variables are trained end-to-end in a tightly coupled fashion, along with minimizing reconstruction loss ensuring that the model learns both temporal dynamics and stochastic variation in a coordinated way.
6.2 Diverse Video Sample Prediction
@@
Except the GRU all the models studied in this work have a latent space that helps them to produce more than one plausible outcome. Although VAE and NF are able to produce diverse samples, they lack the ability to capture temporal diversity, limiting their effectiveness in modeling dynamic, time-evolving structures. Therefore, we have excluded these models when considering diverse video sample prediction.
The results from our simulations using both VRNN and GRU-NF show that each model utilizes its latent space uniquely. Consequently, VRNN demonstrates better performance in terms of MAE, whereas GRU-NF provides better diversity among samples. This distinction suggests different potential applications for each model.
@@
VRNN is a VAE based architecture which relies on approximate likelihood estimation, that can hinder its ability to capture the full complexity of the underlying data distribution. Since GRU-NF is based on NF architecture that uses exact likelihood estimation, it learns a more expressive latent space with flexible transformations, leading to a wider and more complex distribution. The substantial differences between the latent samples lead to a comparatively higher deviation from ground truth at the cost of a wider variety of predictions. This trade-off highlights the strength of exact likelihood methods like NF in capturing the complete probability distribution, which is especially valuable for applications requiring probabilistic forecasting or uncertainty quantification.
@@ 7 Conclusions and Future Work
In this research, we aim to investigate the strengths and limitations of generative time series models for both single and diverse sample prediction in video datasets. We explore a combination of the GRU which is a sequence prediction model with two different generative models, VAE and NF. The first combination, known as VRNN, pairs GRU with VAE, while the second, GRU-NF, integrates GRU with Normalizing Flows.
@@
Our findings indicate that in single video sample prediction, VRNN achieves the best point-forecast fidelity (lowest MAE) and can potentially enable 2x additional bandwidth reduction in existing motion transfer pipelines. It is particularly competitive in datasets that are more ambiguous and multi-modal. In contrast, GRU attains the best performance in settings where the motion is more repeatable or involves subtle changes, thereby suggesting that stable deterministic models can align well with global temporal feature statistics. These findings highlight that the choice of architecture should be guided by the conditional uncertainty of the dataset. In the case of diverse video sample prediction, GRU-NF proves superior in terms of the diversity-fidelity trade-off. While VRNN generates samples closely resembling the ground truth with minimal variation among them, GRU-NF produces more diverse samples without significantly compromising visual quality. This highlights the fundamental distinction between the two approaches: VRNN prioritizes accuracy at the expense of diversity, whereas GRU-NF does the opposite.
@@
The choice between these models depends on the specific application for real-time motion transfer depending on whether the task requires precision or diversity. In scenarios such as video conferencing and remote patient monitoring, where accurate forecasting is critical, VRNN or GRU is preferred depending on dataset characteristics: VRNN is more suitable for higher multi-modality settings e.g., posture changes or irregular patient movement, while GRU can be competitive for more constrained or periodic motion e.g., smooth head movements. Conversely, applications like VR gaming and vision-based anomaly detection in the manufacturing industry require diverse sample generation at future timesteps with high fidelity, making GRU-NF the more suitable option. Our future work will focus on investigation of these trade-offs involved in using various generative time series models for different real-time motion transfer applications on resource constrained edge computing platforms. In addition we plan to enhance FOMM’s lightweight motion transfer pipeline by considering application driven upgrades to its keypoint based motion modeling. This will be done by incorporating more expressive geometric transforms to maintain the required semantic and spatial consistency while explicitly preserving the superior real-time characteristics of FOMM and ease of integration with generative time series models.
Statements and Declarations
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
TH: Methodology, Investigation, Validation, Visualization, Writing – original draft. MABS: Validation, Visualization, Writing –original draft. BJ: Software. XB: Software. SM: Writing – review & editing, SP: Writing – review & editing, IA: Writing – review & editing, SD: Conceptualization, Methodology, Resources, Supervision, Writing – review & editing.
Funding
The author(s) declare that financial support was received for the research of this article. Tasmiah Haque was partially supported for this work by an Intel grant.
Acknowledgments
@@
The authors would like to acknowledge the Pacific Research Platform, NSF Project ACI-1541349, and Larry Smarr (PI, Calit2 at UCSD) for providing the computing infrastructure used in this project. The authors also thank Md Mushfiqur Rahaman for helpful technical discussions and computing support.
Data Availability Statement
The datasets used to conduct this study are available from the corresponding author upon request. Some generated video samples that represent the findings of this research are provided at: https://github.com/Tasmiah1408028/Motion-Transfer-Videos
@@ A. Motion Transfer Backbone (FOMM)
For motion transfer, we utilize the original implementation of FOMM and retain its default configurations for our experiments. For architectural and implementation details, please refer to the supplementary material of [siarohin2019first].
@@ B.1 VRNN
The VRNN model is based on a single-layer GRU with 64-256 hidden units that captures temporal dependencies in 60-dimensional keypoint sequences. It integrates latent-variable modeling using a VAE-like structure. At each timestep, a 15 dimensional latent vector is sampled from a Gaussian distribution using prior and posterior networks, both implemented as feedforward layers with 512 hidden units. The latent code is concatenated with the GRU hidden state and passed through a decoder with a 512-unit hidden layer followed by a linear layer mapping to 60-dimensional keypoints.
@@ B.2 GRU-NF
We use a probabilistic sequential generative model that combines a three-layer GRU network with 512-2048 hidden units per layer, followed by an NF component implemented using 12 RealNVP blocks. The GRU encodes the 60-dimensional input into a hidden representation which is then used to generate a sequence of 60-dimensional latent variables through the RealNVP blocks. Each block contains a pair of scale and translation subnetworks, each composed of five fully connected layers with 2048 hidden units and Tanh/ReLU activations, respectively. These blocks are interleaved with Batch Normalizing layers to stabilize training.